The following fractions represent just three different numbers. Separate them into three groups of equivalent fractions, by changing each one to its simplest form.
12/75

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

12 / 75 = (4 × 3) / (25 × 3)
= 4 / 25

​Thanks